# Heuristical Comparison of Vision Transformers Against Convolutional Neural Networks for Semantic Segmentation on Remote Sensing Imagery

### v1

@@ III Methodology
As shown in Fig. 1, the training lifecycle for our experiment includes five key steps: 1. Dataset Information 2. Data Augmentation, 3. Models, 4. Loss Function 5. Validation Metrics, and 6. Hyperparameters and Training Settings
@@ III-B Data Augmentation
The data augmentation process followed the procedure shown in Algorithm 1. The algorithm first takes anywhere between 6-28% of the image and resizes it to the size of (512, 512) pixels. It then randomly flips the horizontal and vertical axis with a 50% probability and rotates the image anywhere between (0 - 360). Random brightness or contrast is added to the image and is normalized with the mean and standard deviation from the ImageNet dataset. The first step in the process ensures with maximum probability that each of the images we generate from any given image would be either scaled appropriately, have unique features, or both.
@@ III-C Models
We train two models in total. The first one is the UNet based on CNN, and the second one is the fine-tuned MaskFormer ViT.
We use the gradient clipping algorithm by Pascanu et al.[17] to account for exploding gradients in both of our models. Through the heuristic approach, we concluded the maximum value of $threshold$ in algorithm 2 could be set to 3.0.
@@
Each model is different in its fundamental properties and are described below.
@@ III-C1 UNet CNN
Our version of the UNet architecture is a simple adaptation of a generic CNN-based UNet model with four skip connections. The overall architecture of the UNet CNN model is described in Fig. 2. We describe an encoding block and use it throughout the model to do all the feature extraction in both the convolution and transpose convolution layers. The total number of trainable parameters amounts to 42.9M, with the most number of parameters situated in the bottleneck layers. We use gradient accumulation to mimic a batch size of 128 images and use the mixed precision technique with 16 16-bit floats to do the forward pass. These two techniques make the forward process more efficient by skipping backpropagation for n accumulation steps and calculating the predicted mask with just 16-bit precision.
@@ III-C2 MaskFormer ViT
MaskFormer[2] is an algorithm that infuses pixel decoder and transformer decoder with a pre-trained backbone, usually a ResNet[10] block. The image is fed into the backbone, and the resultant is then put forward on the pixel decoder and transformer decoder. The pixel decoder would produce per-pixel embeddings, and the output from the transformer is sent to an MLP. The MLP would produce two results: the N class classification and N mask embeddings. The N mask embeddings are then combined with the per-pixel embeddings, and the output is combined with the N class classification to get the final segmentation. This is represented by Fig 3 extracted from the original paper [2]. They employ their own loss function, which is a combination of the Dice loss function and focal loss function [14]. The number of parameters depends upon the backbone used and ranges all the way from 41M to 212M parameters. For our experiment, we used Facebook’s Swin large architecture, pre-trained on imagenet22k, and fine-tuned for the 15 classes in the iSAID dataset, with ~200M parameters; since most of the parameters are frozen and only the top layer trained, the training efficiency was not affected like the UNet CNN counterpart. The inference time and Floating Operations Per Second (FLOPS) were, however, adversely affected by the larger number of parameters in the model.
@@ III-D Loss Function
$L_{ce}=\beta*A\log(B)+(1-\beta)(1-A)\log(1-B)$ (4)
$L=\lambda_{iou}*L_{iou}+\lambda_{dice}*L_{dice}+\lambda_{ce}*L_{ce}$ (5)
We also did not calculate the loss functions $L_{iou}$ and $L_{dice}$ for the unlabelled class, unlike with $L_{ce}$ because during inference, we need not care about labeling the unlabelled class properly but still the model needs some hint of guidance of the types of patterns it needs to avoid during the training phase, so comes the minimal 0.15 value of $\beta$ in $L_{ce}$. This loss function was only applied to the UNet CNN since the Maskformer model had its own dice and pixel classification loss described in the original paper.
III-E Validation Metrics
$mIoU=\frac{1}{C}\sum_{i=0}^{i=C}\frac{A_{i}\cap B_{i}}{A_{i}\cup B_{i}}$ (6)
$Dice=\frac{1}{C}\sum_{i=0}^{i=C}\frac{2*(A_{i}\cap B_{i})}{|A_{i}|+|B_{i}|}$ (7)
III-F Hyperparameters and Training Settings
Adam[13] was chosen as the optimizer for both models for its capability of fine-tuning the learning rate over time. We trained the models for 40 epochs, each with an initial learning rate of $10^{-3}$ on a virtualized Nvidia A100 GPU with 48G VRAM. The models were validated at the end of each epoch under the Dice score and IoU scores described in (6)(7).
Table I: Model’s details
Model Name # of Parameters FLOPS Inference Time*
UNet CNN 42.9M 460.10 G 0.19s
MaskFormer 200M 232.40 G 0.29s
  • a
    * Inference Time calculated on 6 images
Figure 4: Comparison of Metrics over Epochs
Table II: IoU Scores Comparison by Category
@@ III-F Hyperparameters and Training Settings
    Abbreviations: PL = Plane, BD = Baseball Diamond, BR = Bridge, GTF = Ground Track Field, SV = Small Vehicle, LV = Large Vehicle, SH = Ship, TC = Tennis Court, BC = Basketball Court, ST = Storage Tank, SBF = Soccer Field, RA = Roundabout, HA = Harbor, SP = Swimming Pool, HC = Helicopter. $\dagger$: V=ViT, C=CNN, T=Tiny, S=Small, B=Base
@@ IV Results
Our findings based on the methodology graph shown in Fig 1 are presented in this section. First, we list out the two model’s efficiency information in Table I. During training, most of the time was consumed by the data augmentation technique described in III-B, which took 65 seconds per batch for a batch size of 128. During inference on six images, however, we can see that having more numbers of parameters hurts the efficiency of the MaskFormer model. The UNet CNN had a FLOPS of 460G, whereas the MaskFormer had a FLOPS of 232.4 G. The higher number of FLOPS and lower inference time for the UNet CNN demonstrates its capacity to perform more operations per second even faster, thus resulting in a more efficient model overall.
The MaskFormer ViT model churned better results in terms of mIoU and Dice scores Fig 4. The Dice Score and mIoU for the MaskFormer soared from mid 70% to the range of 80% by the end of the last epoch, whereas for UNet CNN, it started low on around 65% and reached the peak of 81% and stablized at around 78% during training over all classes Fig. 4. The Dice score and mIoU scores have a strong correlation between them, so we further compare the mIoU scores of the model against our references and recent research works. In Table II, recently published and reference works in the topic were selected whose models either have a comparable number of parameters, model architectures, or both as ours to be compared with respect to the mIoU scores and per category IoU scores.
@@
Our UNet CNN model’s metrics on the validation set surpassed the performance of similarly comparable references on the test set, and the MaskFormer, with its greater number of parameters, surpassed the UNet CNN model as well. It can be noted, however, that the mIoU per category of our, or the reference works, were not uniformly distributed, and there’s a high affinity towards objects that tend to be larger in size and more frequent within the dataset. Fig 5 shows sample prediction of randomly selected data against the UNet CNN model and the MaskFormer ViT. The augmentation process shows a high yield of variability in the images, which is capable of generating multiple almost unrecognizable images from a single one. The ground truth was taken from the validation dataset itself, and the UNet and Maskformer’s predictions followed to the right. The segmentation maps in Fig 5, based on the category, show its consistency with the mIoU presented in Table II and the average overall metrics shown in Fig 4.
@@ V Conclusion and Discussions
We successfully show that the introduction of the combined weighted loss function helps the model to make stronger predictions and yield better results. One caveat of the findings could be that we need to find a stronger way to de-segment the testing images into multiple patches of smaller images, which could then be fed into the model and then rearranged later to complete the original images for the testing set, as we learn that simply rescaling does affect the model negatively; in other words, the image loses key patterns and features with simple rescaling in data augmentation. We show the number of parameters required to make a robust remote imagery sensing segmentation model that does not need to be over 50M, given that it can directly influence the FLOPS and inference time. Although the larger training time could be attributed to the data augmentation process, future research could look into streamlining this entire pipeline as a whole. The next steps in the field would entail decoupling the image into multiple smaller fragments and recoupling into the original image to have a robust prediction in images of any shape and size. Nevertheless, we introduced a strong novel combined and weighted loss function to compare UNet CNN with transfer learning-based ViT and show their performance against similar state-of-the-art segmentation models. We show with the use of the new loss function, the CNN model yields comparable results on the metrics with better generalization capacity during inference on unseen data.

### v2

@@ II Literature Review
Xie et al. introduced SegFormer [segformer] that utilize a similar technique to MaskFormer [segformer] for semantic segmentation tasks. Their model performed exceptionally well with high efficiency in terms of performance to efficiency ratio but don’t provide findings in terms of the iSAID datasets in their paper.
@@
From the literary analysis, past research has focused on improving the mIoU of their models (either ViT or CNN) by deploying new architectures, larger models, new loss functions or a combination of either of them. We find research gap on focusing on using a novel loss function that optimizes a relatively smaller CNN UNet based model against fine tuning a relatively larger MaskFormer ViT. We also show that by using our proposed weighted and combined loss function much smaller models can generalize and differentiate on background pixels and yield significantly closer results to models that are much larger than itself. This preserves the model’s ability to retain high mIoU scores while not failing to classify pixels as background at the same time.
@@ III Methodology
Fig. 1 summarizes the training and validation lifecycle of our proposed methodology. The red line segments show the work flow for the training flowchart of the UNet CNN model while the blue line segments show the training/finetuning flowchart for the MaskFormer ViT. As shown in Fig. 1, the training lifecycle for our experiment includes five key steps: A. Dataset Information, B. Data Augmentation, C. Models, D. Loss Function, E. Validation Metrics, and F. Hyperparameters and Training Settings, which are all described below as their own subsections.
@@ III-A Dataset Information
The iSAID dataset is one of the most difficult dataset to generalize a model upon due to its complexities on the ranges of image size and the small number of images. The authors from the dataset [32] acknowledge that any instance segmentation model that performs well on iSAID must have learned robust representations and effective localization capabilities. This would entail that a methodology that performs well on iSAID has a robust foundation to perform well on other segmentation tasks as well.
@@ III-B Data Augmentation
The data augmentation process followed the procedure shown in Algorithm 1. The algorithm first takes anywhere between 6-28% of the image and resizes it to the size of (512, 512) pixels. It then randomly flips the horizontal and vertical axis with a 50% probability and rotates the image anywhere between (0 - 360). Random brightness or contrast is added to the image and is normalized with the mean and standard deviation from the ImageNet dataset. The first step in the process ensures with maximum probability that each of the images we generate from any given image would be either scaled appropriately, have unique features, or both. After the combination of each of the given augmentations, each image can generate hundreds of new images, all of which would have some unique aspects to them.
@@ III-C Models
We trained two models: the first is the UNet based on CNN, and the second is the fine-tuned MaskFormer ViT. For both the models, we use the gradient clipping algorithm by Pascanu et al.[17] to account for exploding gradients. Through the trial and error approach, we concluded the maximum value of $threshold$ in algorithm 2 could be set to 3.0.
@@
Each model is different in its fundamental properties and is described in its own subsection below.
@@ III-C1 UNet CNN
Our version of the UNet architecture is a simple adaptation of a generic CNN-based UNet model with four skip connections. The overall architecture of the UNet CNN model is described in Fig. 2. We describe an encoding block and use it throughout the model to do all the feature extraction in both the convolution and transpose convolution layers. The total number of trainable parameters of Fig.2 amounts to 42.9M, with most of the learnable parameters concentrated in the bottleneck layers. We used gradient accumulation to mimic a batch size of 128 images and employed the mixed precision technique to quantize the forward operations to 16-bit floats. These two techniques make the forward process more efficient by skipping backpropagation for n accumulation steps and calculating the predicted mask with just 16-bit precision.
@@ III-C2 MaskFormer ViT
MaskFormer[2] is an algorithm that infuses pixel decoder and transformer decoder with a pre-trained backbone, usually a ResNet[10] block. The image is forwarded into the backbone and the resultant is passed on to the the pixel decoder and transformer decoder. The pixel decoder would produce per-pixel embeddings, and the transformer’s output after being sent to an MLP gives two results: the N class classification and the N mask embeddings. The N mask embeddings are then combined with the per-pixel embeddings, and the output is combined with the N class classification to get the final segmentation. This is represented by Fig 3 extracted from the original paper [2].
The number of parameters for MaskFormer depends upon the backbone used and ranges from 41M to 212M parameters. For our experiment, we used Facebook’s Swin Large architecture, pre-trained on imagenet22k, and fine-tune it for the 15 classes in the iSAID dataset, resulting to approximately 200M parameters. Since most of the parameters are frozen and only the top layer trained, the training efficiency was faster than the UNet counterpart.
@@ III-D Loss Function
Authors form [2] employ a novel loss function, which is a combination of the Dice loss function and focal loss function [14]. Therefore, we used the same implementation for the fine tuning process. For the UNet, we come up with our proposed weighted loss function which focuses on maximizing mIoU, Dice and minimizing cross-entropy.
@@
Table I: IoU Scores Comparison by Category
@@
    Abbreviations: PL = Plane, BD = Baseball Diamond, BR = Bridge, GTF = Ground Track Field, SV = Small Vehicle, LV = Large Vehicle, SH = Ship, TC = Tennis Court, BC = Basketball Court, ST = Storage Tank, SBF = Soccer Field, RA = Roundabout, HA = Harbor, SP = Swimming Pool, HC = Helicopter. $\dagger$: V=ViT, C=CNN, T=Tiny, S=Small, B=Base.
Figure 4: Comparison of Metrics over Epochs
$L_{ce}=\beta*A\log(B)+(1-\beta)(1-A)\log(1-B)$ (4)
$L=\lambda_{iou}*L_{iou}+\lambda_{dice}*L_{dice}+\lambda_{ce}*L_{ce}$ (5)
We also did not calculate the loss functions $L_{iou}$ and $L_{dice}$ for the unlabeled class, unlike with $L_{ce}$ because during inference we need not care about properly labeling the unlabeled class, but the model still needs some hint of guidance of the types of patterns it needs to avoid during the training phase.
We choose the minimal value of 0.15 for $\beta$ in $L_{ce}$. This loss function in Eq. 5 represents the final form of our proposed loss.
III-E Validation Metrics
$mIoU=\frac{1}{C}\sum_{i=0}^{i=C}\frac{A_{i}\cap B_{i}}{A_{i}\cup B_{i}}$ (6)
$Dice=\frac{1}{C}\sum_{i=0}^{i=C}\frac{2*(A_{i}\cap B_{i})}{|A_{i}|+|B_{i}|}$ (7)
III-F Hyperparameters and Training Settings
Adam[13] was chosen as the optimizer for both models because of its ability to fine-tune the learning rate over time. We trained the models for 40 epochs, each with an initial learning rate of $10^{-3}$ on a Nvidia A40-48Q GPU with 48 GB VRAM. The initial learning rate was chosen so it could easily increase or decrease its value depending on the model’s performance. As Adam updates the learning rate depending on the gradients automatically, all we had to ensure was that the loss was decreasing and the validation metrics were increasing. The models were then validated at the end of each epoch under the Dice score and IoU scores described in (6)(7).
Table II: Models Inference Efficiency Test Results
Model Name # of Parameters FLOPS Inference Time*
UNet CNN 42.9M 460.10 G 0.19s
MaskFormer 200M 232.40 G 0.29s
  • a
    * Inference Time calculated on 6 images
@@ IV Results
Our findings based on the methodology graph shown in Fig 1 are presented in this section.
First, we list out the two model’s efficiency information in Table II. During training, most of the time was consumed by the data augmentation technique described in III-B, which took 65 seconds per batch for a batch size of 128. During inference on six images, however, we can see that having more numbers of parameters hurts the efficiency of the MaskFormer model. This moves us into the question of FLOPS vs Inference Time comparison during inference. From TableII, we notice that UNet CNN has a higher FLOPS and a lower inference time (460.10 G vs 0.19s) as compared to that of MaskFormer (232.40G vs 0.29s). This indicates that on average the UNet model performs better in terms of efficiency by making full utilization of the resources for shorter bursts of time. For where segmentation tasks are mostly used like in UAV, cars, UGV and similar tools, having models which can give output as soon as possible with relatively good accuracy is more important than waiting a few extra seconds to gain a 3-5% increase in the output. This indicates that the UNet CNN model could be a better alternative just by looking at the inference table.
The MaskFormer ViT model, however, churned better results in terms of mIoU and Dice scores Fig 4. The Dice Score and mIoU for the MaskFormer soared from mid 70% to the range of 80% by the end of the last epoch, whereas for UNet CNN, it started low on around 65% and reached the peak of 81% and stablized at around 78% during training over all classes Fig. 4. The Dice score and mIoU scores have a strong correlation between them, so we further compare the mIoU scores of the model against our references and recent research works.
In Table I, recently published and reference works in the topic were selected whose models either have a comparable number of parameters, architectures, or both as ours to be compared with respect to the mIoU scores and per category IoU scores. We have bolded the previous research with comparable or better results in terms of IoU per category in the table. The key advantage here is that the background class is not taken into consideration in any, ours or the previous, of the evaluation. This gives a good advantage to all the models but doesn’t truly show the strong capacity of the weighted combined loss function to generalize and bring about meaningful maps which we further discuss next. Both the models performed exceptionally well, especially in terms of mIoU, but do have some pitfalls in terms of IoU per category where generalization to some classes like small vehicle and large vehicle were better comprehended by previous works. The only problem with this evaluation is we don’t know the exact number of parameters, the inference time and the FLOPS for the models which did better in these specific category in Table I.
@@
Our UNet CNN model’s metrics on the validation set surpassed the performance of similarly comparable references on the test set, and the MaskFormer, with its greater number of parameters, surpassed the UNet CNN model as well. It can be noted, however, that the mIoU per category of our, or the reference works, were not uniformly distributed, and there’s a high affinity towards objects that tend to be larger in size and more frequent within the dataset. Fig 5 shows sample prediction of randomly selected data against the UNet CNN model and the MaskFormer ViT. The augmentation process shows a high yield of variability in the images, which is capable of generating multiple almost unrecognizable images from a single one. The ground truth was taken from the validation dataset itself, and the UNet and Maskformer’s predictions followed to the right. The segmentation maps in Fig 5, based on the category, show its consistency with the mIoU presented in Table I and the average overall metrics shown in Fig 4. This correlation fgvrther validates the experimental design of the combined loss and puts UNet CNN directly against the MaskFormer ViT yielding not only good results in terms of numbers but also in terms of generalization on unseen images, the background class and high fidelity of the segmentation maps.
@@ V Conclusion and Discussions
We successfully show that the introduction of the combined weighted loss function helps the model to make stronger predictions and yield better results. One caveat of the findings could be that we need to find a stronger way to de-segment the testing images into multiple patches of smaller images, which could then be fed into the model and then rearranged later to complete the original images for the testing set, as we learn that simply rescaling does affect the model negatively; in other words, the image loses key patterns and features with simple rescaling in data augmentation. We show the number of parameters required to make a robust remote imagery sensing segmentation model does not need to be over 50M, given that it can directly influence the FLOPS and inference time. We have shown our results in context of previous literature and proved that the usage of the proposed loss does affect the performance model positively and with great significance.
Although the larger training time could be attributed to the data augmentation process, future research could look into streamlining this entire pipeline as a whole. The next steps in the field would entail decoupling the image into multiple smaller fragments and recoupling into the original image to have a robust prediction in images of any shape and size. Nevertheless, we introduced a strong novel combined and weighted loss function to compare UNet CNN with transfer learning-based ViT and show their performance against similar state-of-the-art segmentation models. We show with the use of the new loss function, the CNN model yields comparable results on the metrics with better generalization capacity during inference on unseen data. We also show that the model was better able to make use of the hardware on full capacity yielding the mentioned results while demanding significantly less amount of time.